# An Introduction to Decentralized Training and Execution in Cooperative Multi-Agent Reinforcement Learning

### v3

@@ 2.3.5 Likelihood Q-learning
Each of these ideas can be used in isolation or in combination and they can be used with other decentralized MARL methods to further improve performance. For instance, the combination of Dec-HDRQN (see 2.3.3) with TDL was shown to outperform Dec-HDRQN and forms of lenient Q-learning. Incorporating risk can further improve performance.
3 Decentralized policy gradient methods
Single-agent policy gradient and actor-critic methods can similarly be extended to the decentralized training multi-agent case by allowing each agent to learn separately. These approaches can scale to larger action spaces and have stronger (local) convergence guarantees than the value-based counterparts above.
Policy gradient methods can be used with continuous-action or stochastic policies. For all approaches, we will assume that we have a stochastic policy for each agent parameterized by ${\psi}_{i}$, where ${\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})$ represents the probability agent $i$ will choose action ${a}_{i}$ given the history ${h}_{i}$ and parameters, ${\psi}_{i}$, $\Pr({a}_{i}|{h}_{i},{\psi}_{i})$. This is in contrast to the value-based methods in Section 2, where deterministic policies were generated based on the learned value functions. Like the value-based methods, the decentralized policy gradient methods also assume agents observe their own action, ${a}_{i}$, the resulting observation, ${o}_{i}$, and the joint reward, $r$ at each time step.
3.1 Decentralized REINFORCE
REINFORCE (Williams, 1992) is one of the oldest (and simplest) policy gradient methods. The basic idea is to estimate the value of the policy using Monte Carlo rollouts and then update the policy using gradient ascent. REINFORCE was extended to the cooperative multi-agent case by Peshkin et al. (2000). Importantly, Peshkin et al. (2000) also showed that the joint gradient can be decomposed into decentralized gradients. That is, decentralized gradient ascent will be equivalent to joint gradient ascent. Like the value-based algorithms, ‘independent’ means each agent is learning independently but the data is generated by all agents from the given policies and all agents make updates at the same time (i.e., concurrent learning). Convergence to a local optimum is guaranteed when agents are learning concurrently because they will apply the same algorithm to the same data at synchronized time steps. This is in contrast to the value-based methods, which do not have strong convergence guarantees. These convergence guarantees extend to many other policy gradient methods (as discussed below).
@@ 3.1 Decentralized REINFORCE
0:  Individual actor models ${\pi}_{i}({a}_{i};{h}_{i})$, parameterized by ${\psi}_{i}$
1:  set $\alpha$
@@
Independent REINFORCE isn’t widely used but serves as a basis for the later policy gradient approaches as well as theoretical guarantees.
3.2 Independent actor critic (IAC)
Since REINFORCE is a Monte Carlo method, it must wait until the end of an episode to update. Furthermore, it will typically be less sample efficient than methods that learn a value function. Actor-critic methods have been developed to combat these issues by learning a policy representation (the actor) as well as a value function (the critic) (Sutton and Barto, 2018). Actor-critic methods can use the critic to evaluate the actor without waiting until the end of the episode and TD-learning to update the critic more efficiently.
Independent actor-critic (IAC) was first introduced by Foerster et al. (2018) and it can be thought of as a combination of IQL (or IDRQN) and decentralized REINFORCE. Each agent learns a policy, ${\pi}_{i}({a}_{i};{h}_{i})$, as well as a value function such as ${Q}_{i}({h}_{i},{a}_{i})$. In the simplest case of IAC (where a Q-value is used for the critic), the gradient associated with the policy parameters ${\psi}_{i}$ can be derived as
$\nabla_{\!{\psi}_{i}}J=(1-\gamma)E_{\mathbf{h}{},\mathbf{a}\sim\rho(\mathbf{h}{},\mathbf{a})}\left[{Q}_{i}({h}_{i},{a}_{i})\nabla_{\!{\psi}_{i}}\log{\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})\right]\,,$ (8)
assuming a stochastic policy for each agent $i$ that is parameterized by ${\psi}_{i}$, ${\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})$. Policies are sampled according to the on-policy (discounted) visitation probability $\rho$ (which is a joint visitation probability since concurrent sampling is assumed). This gradient is very similar to the REINFORCE one but uses a Q-value rather than the Monte Carlo return (and is shown here with the expectation rather than just a sample).
@@ 3.2 Independent actor critic (IAC)
$\mathcal{L}({\theta}_{i})=\mathbb{E}_{<{h}_{i},{a}_{i},r,{o}_{i}>\sim\mathcal{D}}\Big{[}\big{(}y-{Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})\big{)}^{2}\Big{]}\text{, where\,\, }y=r+\gamma{Q}_{i}^{{\theta}_{i}}({h}_{i}^{\prime},{a}_{i}^{\prime})$ (9)
where ${h}_{i}^{\prime}={h}_{i}{a}_{i}{o}_{i}$ and ${a}_{i}^{\prime}$ is the action sampled at the next time step using ${\pi}_{i}^{{\psi}_{i}}({a}_{i}^{\prime}|{h}_{i}^{\prime})$ since it is an on-policy estimate (rather than the max action used in DQN). ${Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})$ represents the Q-function estimate for agent $i$ using parameters ${\theta}_{i}$ (just like in Section 2).
Pseudocode for IAC is given in Algorithm 4. Because the value function is no longer being used to select actions but rather just to evaluate the policy, it is common to instead learn ${V}_{i}({h}_{i})$. And because a model of $V$ is learned, we denote it $\hat{V}$. Like the REINFORCE case, the algorithm loops for some number of episodes, choosing actions from the stochastic policy and seeing the joint reward and local observation at each step. Now, during the episode, the TD error is calculated and used to update the actor and the critic using the associated gradients described above. Since V-values are learned instead of Q-values, the TD error is calculated with ${r}_{t}+\gamma\hat{V}_{i}(h_{i,t+1})-\hat{V}_{i}(h_{i,t})$ rather than ${Q}_{i}$. As a result, a value-based version of Equation 8 is used to update the actors. Also, a value-based version of Equation 9 is recovered by using $y=r_{t}+\gamma\hat{V}_{i}(h_{i,t+1})$ and replacing ${Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})$ with $\hat{V}_{i}(h_{i,t})$.
@@
As noted in the value-based (DRQN) case, using an RNN may merge histories of different lengths into the same representation, resulting in incorrectly timed value functions in the finite-horizon case. Like the other algorithms, IAC can be adapted to the episodic or infinite-horizon case by including terminal states or removing the loop over episodes. IAC is a simple baseline that is often used and serves as the basis of many other decentralized MARL actor-critic methods.
3.3 Other decentralized policy gradient methods
Any single-agent method can be extended to the decentralized MARL case. Typically, it just entails including recurrent layers to deal with partial observability and then training agents concurrently to generate joint data and synchronize updates. One notable example is IPPO (de Witt et al., 2020; Yu et al., 2022). IPPO is the extension of single-agent PPO (Schulman et al., 2017) to the multi-agent case. Forms of IPPO have been shown to outperform many state-of-the-art MARL methods. It is worth noting that the standard implementations of IPPO (and the impressive results) use parameter sharing (as discussed below) so they are not technically decentralized but decentralized implementations can still perform well (Yu et al., 2022).
4 Other topics
@@
Parameter (or weight) sharing, where agents share the same network for estimating the value function in value-based methods or policies (and value functions) in policy gradient methods, is common in cooperative MARL. The data from each agent can be used to update a single value network for an algorithm such as DRQN. Agents can still perform differently due to observing different histories and agent indices can be added to increase specialization (Gupta et al., 2017). Since parameter sharing requires agents to share networks during training, it isn’t strictly DTE as it couldn’t be used for online training in a decentralized fashion. Nevertheless, any of the algorithms in this text can be extended to use parameter sharing. The resulting algorithms would be a form of CTDE as the shared network would represent a centralized component. Nevertheless, these parameter-sharing implementations may be more scalable than other forms of CTDE, are often simple to implement, and can perform well (Gupta et al., 2017; Foerster et al., 2018; Yu et al., 2022).
@@
Parameter sharing is one way to exploit a centralized training phase but many others are possible (e.g., sharing a centralized critic (Lowe et al., 2017; Foerster et al., 2018) or learning individual value functions from a joint one (Sunehag et al., 2017; Rashid et al., 2018; Wang et al., 2021)). While CTDE methods are (by far) the most popular form of MARL, they do not always perform the best. In fact, since DTE methods typically make the concurrent learning assumption, they are actually quite close to CTDE methods. This is shown explicitly in the policy gradient case as the gradient of the joint update is the same as the decentralized gradient (as described above in Section 3.1)(Peshkin et al., 2000). This phenomenon has been also been studied theoretically and empirically with modern actor-critic methods (Lyu et al., 2021, 2023). It turns out that while centralized critic actor-critic methods are often assumed to be better than DTE actor-critic methods, they are theoretically the same (with mild assumptions) and often empirically similar (and sometimes worse). Too much centralized information can sometimes be overwhelming, harming scalability of centralized-critic methods (Yu et al., 2022; Lyu et al., 2023). Furthermore, some CTDE methods, such as using a state-based critic, are fundamentally unsound and can often hurt performance in partially observable domains (Lyu et al., 2022, 2023). Of course, there are many different ways of performing centralized training for decentralized execution, and studying the differences and similarities between DTE and CTDE variants of algorithms seems like a promising direction for understanding current methods and developing improved approaches for both cases.

### v4

@@ 3.2.3 Deep extensions, issues, and fixes
Each of these ideas can be used in isolation or in combination and they can be used with other decentralized MARL methods to further improve performance. For instance, the combination of Dec-HDRQN (see 3.2.3) with TDL was shown to outperform Dec-HDRQN and forms of lenient Q-learning. Incorporating risk can further improve performance.
3.3 Decentralized policy gradient methods
Single-agent policy gradient and actor-critic methods can similarly be extended to the decentralized training multi-agent case by allowing each agent to learn separately. These approaches can scale to larger action spaces and have stronger (local) convergence guarantees than the value-based counterparts above. Background on single-agent policy gradient methods is in Section 1.3.2.
Policy gradient methods can be used with continuous-action or stochastic policies. For all approaches, I will assume that we have a stochastic policy for each agent parameterized by ${\psi}_{i}$, where ${\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})$ represents the probability agent $i$ will choose action ${a}_{i}$ given the history ${h}_{i}$ and parameters, ${\psi}_{i}$, $\Pr({a}_{i}|{h}_{i},{\psi}_{i})$. This is in contrast to the value-based methods in Section 3.2, where deterministic policies were generated based on the learned value functions.
3.3.1 Decentralized REINFORCE
As mentioned in Section 1.3.2, REINFORCE (Williams, 1992) is one of the oldest (and simplest) policy gradient methods. The basic idea is to estimate the value of the policy using Monte Carlo rollouts and then update the policy using gradient ascent. REINFORCE was extended to the cooperative multi-agent case by Peshkin et al. (2000). Importantly, Peshkin et al. (2000) also showed that the joint gradient can be decomposed into decentralized gradients. That is, decentralized gradient ascent will be equivalent to joint gradient ascent. Convergence to a local optimum is guaranteed when agents are learning concurrently (each agent is learning on the data that is generated by all agents and all agents make updates at the same time) because they will apply the same algorithm to the same data at synchronized time steps. This convergence result is in contrast to the value-based methods, which do not have strong convergence guarantees. These convergence guarantees extend to many other policy gradient methods (as discussed below).
@@ 3.3.1 Decentralized REINFORCE
0:  Individual actor models ${\pi}_{i}({a}_{i}|{h}_{i})$, parameterized by ${\psi}_{i}$
1:  set $\alpha$ (learning rate)
@@
Decentralized REINFORCE isn’t widely used but serves as a basis for the later policy gradient approaches as well as theoretical guarantees.
3.3.2 Independent actor critic (IAC)
Since REINFORCE is a Monte Carlo method, it must wait until the end of an episode to update. Furthermore, it will typically be less sample efficient than methods that learn a value function. As mentioned in Section 1.3.2, actor-critic methods have been developed to combat these issues by learning a policy representation (the actor) as well as a value function (the critic) (Sutton and Barto, 2018). Actor-critic methods can use the critic to evaluate the actor without waiting until the end of the episode and TD-learning to update the critic more efficiently.
Independent actor-critic (IAC) was first introduced by Foerster et al. (2018b) and it can be thought of as a combination of IQL (or IDRQN) and decentralized REINFORCE. Each agent learns a policy, ${\pi}_{i}({a}_{i}|{h}_{i})$, as well as a value function such as ${Q}_{i}({h}_{i},{a}_{i})$. In the simplest case of IAC (where a Q-value is used for the critic), the gradient associated with the policy parameters ${\psi}_{i}$ can be derived as
$\nabla_{\!{\psi}_{i}}J=\mathbb{E}_{\mathbf{h}{},\mathbf{a}\sim\mathcal{D}}\left[{Q}_{i}({h}_{i},{a}_{i})\nabla_{\!{\psi}_{i}}\log{\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})\right]\,,$ (3.6)
assuming a stochastic policy for each agent $i$ that is parameterized by ${\psi}_{i}$, ${\pi}_{i}^{{\psi}_{i}}({a}_{i}|{h}_{i})$. Policies are sampled according to the on-policy (discounted) visitation probability (which is a joint visitation probability since concurrent sampling is assumed). This gradient is very similar to the REINFORCE one but uses a Q-value rather than the Monte Carlo return (and is shown here with the expectation rather than just a sample).
@@ 3.3.2 Independent actor critic (IAC)
$\mathcal{L}({\theta}_{i})=\mathbb{E}_{<{h}_{i},{a}_{i},r,{o}_{i}>\sim\mathcal{D}}\Big{[}\big{(}y-{Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})\big{)}^{2}\Big{]}\text{, where\,\, }y=r+\gamma{Q}_{i}^{{\theta}_{i}}({h}_{i}^{\prime},{a}_{i}^{\prime})$ (3.7)
where ${h}_{i}^{\prime}={h}_{i}{a}_{i}{o}_{i}$ and ${a}_{i}^{\prime}$ is the action sampled at the next time step using ${\pi}_{i}^{{\psi}_{i}}({a}_{i}^{\prime}|{h}_{i}^{\prime})$ since it is an on-policy estimate (rather than the max action used in DQN). ${Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})$ represents the Q-function estimate for agent $i$ using parameters ${\theta}_{i}$ (just like in Section 3.2).
Pseudocode for IAC is given in Algorithm 4. Because the value function is no longer being used to select actions but rather just to evaluate the policy, it is common to instead learn ${V}_{i}({h}_{i})$. And because a model of $V$ is learned, I denote it $\hat{V}$. Like the REINFORCE case, the algorithm loops for some number of episodes, choosing actions from the stochastic policy and seeing the joint reward and local observation at each step. Now, during the episode, the TD error is calculated and used to update the actor and the critic using the associated gradients described above. Since V-values are learned instead of Q-values, the TD error is calculated with ${r}_{t}+\gamma\hat{V}_{i}(h_{i,t+1})-\hat{V}_{i}(h_{i,t})$ rather than ${Q}_{i}$. As a result, a value-based version of Equation 3.6 is used to update the actors. A value-based version of Equation 3.7 is recovered by using $y=r_{t}+\gamma\hat{V}_{i}(h_{i,t+1})$ and replacing ${Q}_{i}^{{\theta}_{i}}({h}_{i},{a}_{i})$ with $\hat{V}_{i}(h_{i,t})$.
@@
Like the other algorithms, IAC can be adapted to the episodic or infinite-horizon case by including terminal states or removing the loop over episodes. IAC is a simple baseline that is often used and serves as the basis of many other decentralized MARL actor-critic methods.
3.3.3 Other decentralized policy gradient methods
Any single-agent method can be extended to the decentralized MARL case. Typically, it just entails including recurrent layers to deal with partial observability and then training agents concurrently to generate joint data and synchronize updates. One notable example is IPPO (de Witt et al., 2020; Yu et al., 2022). IPPO is the extension of single-agent PPO (Schulman et al., 2017) to the multi-agent case. Forms of IPPO have been shown to outperform many state-of-the-art MARL methods. It is worth noting that the standard implementations of IPPO (and the impressive results) use parameter sharing (as discussed below) so they are not technically decentralized but decentralized implementations can still perform well (Yu et al., 2022). I discuss IPPO (and MAPPO) in more detail in Section 4.3.5 and discuss parameter sharing below.
3.4 Other topics
@@
Parameter (or weight) sharing, where agents share the same network for estimating the value function in value-based methods or policies (and value functions) in policy gradient methods, is common in cooperative MARL. The data from each agent can be used to update a single value network for an algorithm such as DRQN. Agents can still perform differently due to observing different histories and agent indices can be added to increase specialization (Gupta et al., 2017; Foerster et al., 2016). While parameter sharing is typically used with homogeneous agents (i.e., those with the same action and observation spaces), it can also be used with heterogeneous agents (Terry et al., 2020). Since parameter sharing requires agents to share networks during training, it isn’t strictly DTE as it couldn’t be used for online training in a decentralized fashion. Nevertheless, any of the algorithms in this chapter can be extended to use parameter sharing. The resulting algorithms would be a form of CTDE as the shared network would represent a centralized component. Nevertheless, these parameter-sharing implementations may be more scalable than other forms of CTDE, are often simple to implement, and can perform well (Gupta et al., 2017; Foerster et al., 2018b; Yu et al., 2022).
@@
While CTDE methods are (by far) the most popular form of MARL, they do not always perform the best. In fact, since DTE methods typically make the concurrent learning assumption, they are actually quite close to CTDE methods. This is shown explicitly in the policy gradient case as the gradient of the joint update is the same as the decentralized gradient (as described above in Section 3.3.1)(Peshkin et al., 2000). This phenomenon has been also been studied theoretically and empirically with modern actor-critic methods (Lyu et al., 2021, 2023). It turns out that while centralized critic actor-critic methods are often assumed to be better than DTE actor-critic methods, they are theoretically the same (with mild assumptions) and often empirically similar (and sometimes worse). Too much centralized information can sometimes be overwhelming, harming scalability of centralized-critic methods (Yu et al., 2022; Lyu et al., 2023). Furthermore, some CTDE methods, such as using a state-based critic, are fundamentally unsound and can often hurt performance in partially observable domains (Lyu et al., 2022, 2023). More details about these issues can be found in Sections 4.3.7 and 4.3.6. Of course, there are many different ways of performing centralized training for decentralized execution, and studying the differences and similarities between DTE and CTDE variants of algorithms seems like a promising direction for understanding current methods and developing improved approaches for both cases.